# Adaptive refinement for eigenvalue problems based on an associated source problem

### v1

@@ 1. Introduction
Adaptive finite element schemes for eigenvalue computations have a long history, with contributions typically mirroring (and often following soon after) their counterparts for source problems. In the case of computing clusters of eigenvalues, where repeated eigenvalues (or merely tightly clustered eigenvalues) may be present, together with bases for their corresponding invariant subspaces is more a challenging task, and contributions in this direction is relatively recent in comparison with those that concern a few simple eigenvalues near the bottom of the spectrum. We do not attempt a comprehensive survey of the extensive literature on the topic, and instead provide some references that more specifically concern the approximation of several, possibly clustered, eigenvalues, together with a basis of the corresponding invariant subspace [23, 5, 14, 7, 8, 16, 20, 28]. A common theme of nearly all of these contributions it that, whatever strategy one ultimately chooses for adaptively enriching the finite element space used to approximate the eigenpairs, the resulting finite element space should be well-suited to approximate generic functions in the entire invariant subspace. An exception to this pattern is the work [20], in which multiple meshes/FEM spaces are employed.
The generic template for adaptive finite element procedures, for source problems or eigenvalue problems, is to repeat the the following steps until some convergence criterion is met:
@@
It is possible to construct examples, particularly in 1D, for which the landscape function $u$ is a polynomial (of low degree) in some subdomain $R$, but for which the eigenvectors $\psi_{j}$ of interest are not polynomial on $R$. For such an example we might rightly expect refinement driven by approximations of $u$ to cease refining within $R$ at some stage, and thereby fail to provide convergent eigenvector approximations. Such situations are highly unlikely in practice, however, and we will provide both theory substantial empirical evidence that “landscape refinement” is a very effective approach for the efficient approximation of (large) clusters of eigenvalues and eigenvectors. In Example 4.5, we illustrate a different situation in which a few eigenvectors within a large cluster are converging more slowly under landscape refinement than the rest, and we indicate how such eigenvectors can be detected, extracted, and dealt with efficiently.
The rest of the paper is structured as follows. In Section 2, we provide theoretical results that link the landscape function $u$ to eigenpairs $(\lambda_{j},\psi_{j})$ in a way that makes the efficient approximation of such eigenvectors on finite element spaces based on meshes generated by landscape refinement more reasonable than it may appear at first glance. Preliminary experiments in this section add to the evidence that landscape refinement might actually be a viable alternative to refinement based on the (approximate) eigenpairs of interest. Corollary 2.2, which concerns the pointwise control of properly scaled eigenvectors by $u$ is, by now, a well-known result (cf. [13, 3, 4, 11]). Our more general result, Theorem 2.1, does not appear to be broadly known—we have not found it in the literature. Apart from the initial proof-of-concept computations in Section 2, our numerical experiments are done using the SIPG variant of $hp$ DG finite element discretizations. These are briefly described in Section 3, together with the error estimators and local error indicators that we use in our experiments. Section 4 contains the bulk of the experiments, in which we illustrate the effectiveness of our approach on a variety of problems making relevant comparisons with more traditional approaches. Details about the algorithms used for these comparisons are given in Section 5, and further enhancements of our proposed algorithm are considered in Section 6. A few concluding remarks are given in Section 7
@@ 2. Basic Properties of the Landscape Function
(5) $u=\sum_{n\in\mathbb{N}}c_{n}\,\psi_{n}\quad,\quad c_{n}=\left(\int_{\Omega}\psi_{n}\,dx\right)/\lambda_{n}~{}.$
@@
The discussion above was intended to establish some connections between the eigenvectors of $\mathcal{L}$ and the landscape function $u$, as a first indication that adaptive refinement that is aimed at approximating the landscape function might, in fact, yield finite element spaces in which the eigenvectors are also well approximated. We now provide a first illustration of this approach and its performance on a model problem for which there is no eigenvector localization, but for which adaptive refinement is advisable. In Section 4, we provide strong supporting evidence for the “landscape refinement” approach on variety of problems.
@@
Although CR wins in each head-to-head comparison of error, the errors are comparable and decrease at the correct rate in both cases, and LR is computationally cheaper, so LR might be considered a viable alternative. In later experiments, in which larger collection of eigenvalues are desired, we will see that LR becomes an even more attractive alternative.
@@
We next consider a cluster of eigenpairs a bit higher in the spectrum. The interval $[50\pi^{2}-10,50^{2}\pi+10]$ contains seven eigenvalues of $\mathcal{L}$; four are simple, but $\lambda=50\pi^{2}$ has a three-dimensional eigenspace
@@
In this case, we consider three different refinement strategies: landscape refinement (LR), eigenvector refinement (CR), and a mixed strategy (MR) that begins with LR and then shifts to CR when the dimension of the finite element space reaches some threshold. The plots in Figure 5 contain the computed eigenvalues in the interval $[50\pi^{2}-10,50^{2}\pi+10]$ for both the LR and CR refinement strategies, starting from the same coarse mesh as before. Early in the adaptive cycle, neither approach has produced a finite element space that is sufficient to resolve the invariant subspace of interest—for the first several refinement cycles, neither approach even computes the correct the number of eigenvalues in the interval interest! In the case of CR, this means that refinement during this phase is being at least partially driven by approximations of that are, at best, related to eigenvectors of $\mathcal{L}$ for eigenvalues outside of $[50\pi^{2}-10,50^{2}\pi+10]$. If CR is generating useful meshes during this phase, it essentially doing so “on accident”. In fact, CR does seem to be generating useful meshes for approximation the invariant subspace of interest, and it does so a bit earlier than LR refinement, based on when we see a discernable pattern of convergence toward the correct eigenvalues; for CR this happens somewhere around 3000 or 4000 DOF, and for LR it is nearer to 8000 DOF.
@@
Table 1. Example 2.5. Errors $e_{j}=|\lambda_{j}-\hat{\lambda}_{j}|$ on the finest mesh for approximating the seven eigenvalues in $[50\pi^{2}-10,50^{2}\pi+10]$ using several adaptive refinement strategies.
@@ 4. Numerical Results
We will compare landscape refinement with versions of eigenvector-based refinement on several examples having different features. We first define of all the quantities used in presenting the results that are collected, starting from the quantities related to the eigenvalue problem:
@@
Figure 7. Example 4.1. Left: Convergence plots of the relative eigenvalue errors $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}$ under landscape refinement and single eigenpair refinement for $j=1,48,90$. Right: Convergence plots of the relative error envelope bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ under landscape refinement two variants of cluster refinement. Also included is the error bound for the landscape function $\eta_{\mathrm{land}}^{2}$ under landscape refinement.
@@
We briefly revisit Example 2.5, but with $hp$ adaptivity using landscape refinement and cluster refinement for the first $M=100$ eigenpairs. As before, we start with initial mesh of $8\times 8$ elements of order $p=2$. The convergence results for landscape refinement and cluster refinement using AdaptSpectrumEigSum are shown in Figure 8, for the envelope for the relative error bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$. The final meshes in each case are given in Figure 9. We see that landscape refinement, though initially providing slightly slower convergence, overtakes cluster refinement at around $140^{2}$ DOFs, and reaches an error tolerance of $10^{-4}$ for the cluster much(!) sooner. The final mesh in the case of landscape refinement is also more appealing, with $h$-refinement targeted near the corners of the domain, whereas the final mesh for cluster refinement is far less coherent.
Figure 8. Example 4.2. Convergence plots of the relative error envelope bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ under landscape refinement two variants of cluster refinement. Also included is the error bound for the landscape function $\eta_{\mathrm{land}}^{2}$ under landscape refinement.

### v2

@@ 1. Introduction
Adaptive finite element schemes for eigenvalue computations have a long history, with contributions typically mirroring (and often following soon after) their counterparts for source problems. In the case of computing clusters of eigenvalues, where repeated eigenvalues (or merely tightly clustered eigenvalues) may be present, together with bases for their corresponding invariant subspaces is more a challenging task, and contributions in this direction are relatively recent in comparison with those that concern a few simple eigenvalues near the bottom of the spectrum. We do not attempt a comprehensive survey of the extensive literature on the topic, and instead provide some references that more specifically concern the approximation of several, possibly clustered, eigenvalues, together with a basis of the corresponding invariant subspace [24, 5, 15, 7, 9, 17, 21, 30]. A common theme of nearly all of these contributions it that, whatever strategy one ultimately chooses for adaptively enriching the finite element space used to approximate the eigenpairs, the resulting finite element space should be well-suited to approximate generic functions in the entire invariant subspace. An exception to this pattern is the work [21], in which multiple meshes/FEM spaces are employed.
The generic template for adaptive finite element procedures, for source problems or eigenvalue problems, is to repeat the following steps until some convergence criterion is met:
@@
As will be seen in Corollary 2.2 and Proposition 2.4, and in the experiments, we expect the landscape function to reliably encode information concerning where eigenvectors are singular or smooth, or where they exhibit other interesting features such as localization. As such, we contend that refinement based on features of the landscape function will typically provide a family of adapted meshes that resolve the eigenvectors of interest at an optimal rate, and will do so more efficiently than the eigenvector cluster approach, particularly in the case of $hp$-adaptivity.
@@
It is possible to construct examples, particularly in 1D, for which the landscape function $u$ is a polynomial (of low degree) in some subdomain $R$, but for which the eigenvectors $\psi_{j}$ of interest are not polynomial on $R$. For such an example we might rightly expect refinement driven by approximations of $u$ to cease refining within $R$ at some stage, and thereby fail to provide convergent eigenvector approximations. Such situations are highly unlikely in practice, however, and we will provide both theory and substantial empirical evidence that “landscape refinement” is a very effective approach for the efficient approximation of (large) clusters of eigenvalues and eigenvectors. Such examples illustrate why a result of the form $\|\psi-\psi_{hp}\|\leq C\|u-u_{hp}\|$ cannot hold in general. However, the preceding paragraph suggests why such a result does not need to hold for the approach to be effective. In particular, although we do compute a posteriori estimates of $\|u-u_{hp}\|$ for some comparisons, these are not used as part of any stopping criterion. In Example 4.5, we illustrate a different situation in which a few eigenvectors within a large cluster are converging more slowly under landscape refinement than the rest, and we indicate what goes wrong in this case, and how to efficiently restore the proper rate of convergence.
The rest of the paper is structured as follows. In Section 2, we provide theoretical results that link the landscape function $u$ to eigenpairs $(\lambda_{j},\psi_{j})$ in a way that makes the efficient approximation of such eigenvectors on finite element spaces based on meshes generated by landscape refinement more reasonable than it may appear at first glance. Preliminary experiments in this section add to the evidence that landscape refinement might actually be a viable alternative to refinement based on the (approximate) eigenpairs of interest. Corollary 2.2, which concerns the pointwise control of properly scaled eigenvectors by $u$ is, by now, a well-known result (cf. [14, 3, 4, 12]). Our more general result, Theorem 2.1, does not appear to be broadly known—we have not found it in the literature. Apart from the initial proof-of-concept computations in Section 2, our numerical experiments are done using the SIPG variant of $hp$ DG finite element discretizations. These are briefly described in Section 3, together with the error estimators and local error indicators that we use in our experiments. Section 4 contains the bulk of the experiments, in which we illustrate the effectiveness of our approach on a variety of problems, making relevant comparisons with more traditional approaches. Details about the algorithms used for these comparisons are given in Section 5, and further enhancements of our proposed algorithm are considered in Section 6. A few concluding remarks are given in Section 7
@@ 2. Basic Properties of the Landscape Function
[…]
@@
The eigenbasis expansion (5) indicates how features of the eigenvectors are encoded in the landscape function $u$. More specifically, this expansion demonstrates why we generally expect the landscape function to identify the locations and strengths of all singularities that may occur in eigenvectors, highlighting the strongest singularities that can occur at each singular point. In particular $u$ will have the same singularities as the ground state eigenvector $\psi_{1}$, which often exhibits the strongest singularities that can occur for any eigenvector. The causes and strengths of singularities for the types of problems we consider are well-known (cf.[8, 22, 23, 27, 28, 34]). More specifically, these authors identify the kinds of singularities that can occur in solutions source problems $\mathcal{L}v=f$ under reasonable regularity assumptions on $f$. This theory clearly applies to the case of eigenvalue problems, by taking $f=\lambda v$, so it provides a rich understanding of the sorts of singularities eigenvectors may exhibit. Assuming that $u$ does encode the possible singularities of eigenvectors, and that adaptive mesh refinement for $u$ is done in such a way as to generate meshes/spaces that yield optimal order convergence of approximations of $u$, it is unsurprising that computed eigenvectors also converge at the optimal order, at least asymptotically. This is precisely the behavior we observe in nearly all of our experiments—Example 4.5, briefly discussed below, offers a counter-example.
It is possible to construct problems for which the source term $f=1$ is orthogonal to all eigenvectors having a given singular behavior, in which case the landscape function will fail to identify the singularity. This is almost certainly what is happening in Example 4.5, for which we observe that the (empirically) most singular eigenvectors converge more slowly under landscape refinement than the rest. In such a case, one might naturally consider using a different source problem $\mathcal{L}u_{f}=f$ to drive adaptivity, possibly in conjunction with the landscape function. For example, if it appears that $1$ is orthogonal to all eigenvectors having a given singularity, then choosing a simple non-trivial $f$ that is orthogonal to $1$ may identify, via $u_{f}$, the correct singular behavior of the eigenvectors that was missed by $u$. In fact, one expects a “randomly chosen” $f$ to do the job. Understood in that light, one might wonder why we propose $f=1$ as the default choice. Theorem 2.1, Corollary 2.2 and Example 2.3, provide justification that this is a very sensible choice, and our experiments provide further support. Furthermore, the body of work exemplified by the aforementioned contributions [14, 3, 4, 12] demonstrates the perhaps surprising amount of information $u$ reveals about the spectrum of Schrödinger operators. Additionally, the choice $f=1$ does not introduce any particular biases toward certain features that may not be present in most eigenvectors.
We are primarily interested in employing adaptive refinement based on approximations of the landscape function when computing moderate or large collections of eigenpairs. In that setting, it is probable that some (many) of the eigenvectors of interest will have the strongest singularities that are possible for the given domain and differential operator, and our adaptive scheme is expected to yield convergence of collective measures of error that are not only asymptotically optimal, but also quite reasonable in the pre-asymptotic regime. Of course, our one-mesh-fits-all approach will be overly aggressive in its refinement toward singular points in terms of the approximation of smooth(er) eigenvectors, but we do not see this as a disqualifying drawback. At any rate, refinement based on clusters of computed eigenvectors that contain some smooth and some singular eigenvectors would suffer the same drawback, and do so at greater computational cost.
The discussion above was intended to establish some connections between the eigenvectors of $\mathcal{L}$ and the landscape function $u$, as a first indication that adaptive refinement that is aimed at approximating the landscape function might, in fact, yield finite element spaces in which the eigenvectors are also well approximated. We now provide a first illustration of this approach and its performance on a model problem for which there is no eigenvector localization, but for which adaptive refinement is advisable. In Section 4, we provide strong supporting evidence for the “landscape refinement” approach on a variety of problems.
@@
Although CR wins in each head-to-head comparison of error, the errors are comparable and decrease at the correct rate in both cases, and LR is computationally cheaper, so LR might be considered a viable alternative. In later experiments, in which a larger collection of eigenvalues is desired, we will see that LR becomes an even more attractive alternative.
@@
We next consider a cluster of eigenpairs a bit higher in the spectrum. The interval $[50\pi^{2}-10,{50\pi^{2}}+10]$ contains seven eigenvalues of $\mathcal{L}$; four are simple, but $\lambda=50\pi^{2}$ has a three-dimensional eigenspace
@@
In this case, we consider three different refinement strategies: landscape refinement (LR), eigenvector refinement (CR), and a mixed strategy (MR) that begins with LR and then shifts to CR when the dimension of the finite element space reaches some threshold. The plots in Figure 5 contain the computed eigenvalues in the interval $[50\pi^{2}-10,{50\pi^{2}}+10]$ for both the LR and CR refinement strategies, starting from the same coarse mesh as before. Early in the adaptive cycle, neither approach has produced a finite element space that is sufficient to resolve the invariant subspace of interest—for the first several refinement cycles, neither approach even computes the correct number of eigenvalues in the interval interest! In the case of CR, this means that refinement during this phase is being at least partially driven by approximations of that are, at best, related to eigenvectors of $\mathcal{L}$ for eigenvalues outside of $[50\pi^{2}-10,{50\pi^{2}}+10]$. If CR is generating useful meshes during this phase, it essentially doing so “on accident”. In fact, CR does seem to be generating useful meshes for approximating the invariant subspace of interest, and it does so a bit earlier than LR refinement, based on when we see a discernable pattern of convergence toward the correct eigenvalues; for CR this happens somewhere around 3000 or 4000 DOF, and for LR it is nearer to 8000 DOF.
@@
Table 1. Example 2.5. Errors $e_{j}=|\lambda_{j}-\hat{\lambda}_{j}|$ on the finest mesh for approximating the seven eigenvalues in $[50\pi^{2}-10,{50\pi^{2}}+10]$ using several adaptive refinement strategies.
@@ 4. Numerical Results
We will compare landscape refinement with versions of eigenvector-based refinement on several examples having different features. We first define all the quantities used in presenting the results that are collected, starting from the quantities related to the eigenvalue problem:
@@
Figure 7. Example 4.1. Left: Convergence plots of the relative eigenvalue errors $\eta_{\mathrm{eig},j}^{2}/\lambda_{j,h}$ under landscape refinement and single eigenpair refinement for $j=1,48,90$. Right: Convergence plots of the relative error envelope bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ under landscape refinement and two variants of cluster refinement. Also included is the error bound for the landscape function $\eta_{\mathrm{land}}^{2}$ under landscape refinement.
@@
We briefly revisit Example 2.5, but with $hp$ adaptivity using landscape refinement and cluster refinement for the first $M=100$ eigenpairs. As before, we start with an initial mesh of $8\times 8$ elements of order $p=2$. The convergence results for landscape refinement and cluster refinement using AdaptSpectrumEigSum are shown in Figure 8, for the envelope for the relative error bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$. The final meshes in each case are given in Figure 9. We see that landscape refinement, though initially providing slightly slower convergence, overtakes cluster refinement at around $140^{2}$ DOFs, and reaches an error tolerance of $10^{-4}$ for the cluster much(!) sooner. The final mesh in the case of landscape refinement is also more appealing, with $h$-refinement targeted near the corners of the domain, whereas the final mesh for cluster refinement is far less coherent.
Figure 8. Example 4.2. Convergence plots of the relative error envelope bound $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ under landscape refinement and two variants of cluster refinement. Also included is the error bound for the landscape function $\eta_{\mathrm{land}}^{2}$ under landscape refinement.
@@
In the discussion preceding Example 2.5 we indicated that is it possible that the source term $f=1$ used for our landscape function could be orthogonal to all eigenvectors of a certain type, e.g. all eigenvectors possessing the strongest point singularity, though that is expected to be unlikely in practice. In that same discussion, we briefly mentioned the possibility of using a different source term $f$, and driving adaptivity based on the associated $u_{f}$. In our discussion for the current example, we suggested that this orthogonality is probably what is causing the slower convergence of a few eigenvectors here. The results presented in Figure 13 provide strong support of this claim, and also provide a simple potential alternative to flagging slow-converging eigenvectors for special treatment. We simply use a different source term $f$, as indicated above.
Our first variation of the original problem is to move the singular point from $(1/2,1/2)$ to $(3/4,3/4)$ by modifying the diffusion coefficient $A$. In this variant, $A=1$ on $(0,0.75]^{2}$ and $[0.75,1)^{2}$, and $\beta^{2}=10$ elsewhere. This shifting of the singularity results (empirically) in $f=1$ no longer being orthogonal to the eigenvectors of the new operator. The top row in Figure 13 provides a comparison between the behavior under landscape refinement for the original problem and this variant, and we see that the landscape function for the variant appears to encode the worst singular behavior in the eigenvectors. More specifically, the stagnating convergence observed in the original problem is broken at around $100^{2}$ DOFs, and there are no longer any slowly converging eigenvectors.
Our second variation of the original problem keeps the singular point at $(1/2,1/2)$, but uses the source term $f=1-3x$ to drive adaptivity based on $u_{f}$. We highlight that there is nothing particularly special about this choice of $f$ beyond it being a low-degree polynomial that happens to be non-orthogonal to the most singular eigenvectors. The bottom plot in Figure 13 compares the use of $f=1-3x$ and $f=1$. This again corroborates our assertion that the culprit behind the slow convergence of some eigenvectors in the original example was the orthogonality of $f=1$ and those vectors, which led to $u$ failing to encode those singularities. It is also clear that a modest change to $f$ completely fixed this issue. The effectiveness of this alternate approach hints at a modification of our initial strategy of doing adaptive refinement based on a single source problem: choose a small number of fixed sources, e.g. $\{1,x,y\}$, and drive adaptivity based on the the associated solutions. It becomes increasingly unlikely that (singular) eigenvectors are orthogonal to even a small set of source terms. We do not pursue this idea further here.
Figure 13. Example 4.5. Top left: Convergence plots of the square of the error bound for the landscape solution $\eta_{\mathrm{land}}^{2}$ and the envelope for the relative error bounds $\eta_{\mathrm{max},\mathrm{rel}}^{2}$ computed using landscape refinement with singular point at $(1/2,1/2)$ and at $(3/4,3/4)$. Top right: Convergence rates of relative eigenvalue error estimates under landscape refinement for several eigenvalues within the cluster when the singular point is at $(3/4,3/4)$. Bottom: Refinement driven by $u_{f}$ for the source term $f=1-3x$ compared with that of $u$ ($f=1$), with singular point at $(1/2,1/2)$.
Example 4.6 (Perforated Domain).
The family of domains $\Omega_{m}$ obtained by deleting $m^{2}$ squares having edge lengths $H=1/(2m+1)$ from the unit square, as pictured in Figure 14 for $m=3$, provides a setting in which some eigenvectors have many singularities, but other eigenvectors are analytic and known explicitly. More specifically, if we consider the Dirichlet Laplacian on $\Omega_{m}$, many eigenvectors, including the ground state, will have $4m^{2}$ point singularities of type $r^{2/3}$, one for each of the non-convex corners in the domain. The landscape function $u$ will also have these same point singularities. However, there are many analytic eigenvectors as well. For example, $\psi=\sin(j\pi x/H)\sin(k\pi y/H)$ is an eigenvector with corresponding eigenvalue $\lambda=(j\pi/H)^{2}+(k\pi/H)^{2}$ for $j,k\in\mathbb{N}$. The smallest of these, which we expect to belong to the first analytic eigenvector, is $\lambda=2(\pi/H)^{2}$.
When $m=3$, the first 100 eigenvalues are in the interval $[301.46,2070.55]$, and this interval only includes one of analytic eigenvectors described above, namely $\lambda=2(7\pi)^{2}\approx 967.22123$. It is 41st in the spectrum, i.e. $\lambda=\lambda_{41}$. The 41st computed eigenvector approximates $\sin(7\pi x)\sin(7\pi y)$ (up to scaling) reasonably well even on the coarsest space, which consists of quadratic elements on the mesh shown in Figure 14. Our empirical analyticity measures indicate that this is the only smooth eigenvector among the first 100, which aligns with our expectations. All other eigenvectors have analyticity measures similar to that of $u$, which suggests that each has a point singularity at at least one of the non-convex corners. This again agrees with our expectations for this problem.
Because of the large number of singularities for $u$, one might naturally wonder if landscape refinement yields sub-optimal convergence for smooth eigenvectors such as $\psi_{41}$, because it would seem to concentrate much of its “effort” on resolving singularities that are not present in such eigenvectors. As seen in Figure 14, this intuition is correct. Refinement based on the computed approximation of $\psi_{41}$ first achieves eigenvalue errors below $10^{-4}$ at around 40K DOFs, whereas this error is not achieved for landscape refinement until around 120K DOFs. We note, however, that the best observed convergence rates for both approaches are very similar—the landscape approach just takes longer to achieve that rate. This example provides a more dramatic instance of behavior we have seen in Figures 7 (left) and 10 (top right), namely that convergence of a single eigenpair under refinement that is (ideally) tailored to that specific eigenpair can outperform a more generic approach such as landscape refinement. It is perhaps more noteworthy that landscape refinement sometimes “wins” against single eigenvector refinement in this setting, and these same figures show that it can. Regardless, we primarily promote landscape refinement as an efficient means of driving down collective measures of error (e.g. max error) for a cluster of eigenvalues, and such clusters will generically contain eigenvectors having singularities like those of $u$. If one is willing to compute local analyticity measures, which is not uncommon in $hp$-adaptive algorithms, any eigenvectors that are flagged as being (much) smoother than $u$ could be extracted from a cluster for different treatment, in a similar fashion as we proposed in Example 4.5 for eigenvectors that are more singular than $u$.
Figure 14. Example 4.6. Top row: The perforated unit square with $m^{2}$ square holes ($m=3$), and the initial mesh. Bottom row: Eigenvalue convergence under landscape refinement for several eigenvalues, including $\lambda_{41}$, which has the only analytic eigenvector among the first 100. Comparison of convergence for $\lambda_{41}$ under landscape refinement and single eigenvector refinement based on the computed approximation of $\psi_{41}$.